OpenStudy (anonymous):

Fill in the reason that justifies the step by step. Given: QS = 42 Prove: x = 13 will medal and fan...

OpenStudy (anonymous):

QS= 42 QR+RS= QS (x+3) + 2x=42 3x+3=42 3x=39 x=13

jhonyy9 (jhonyy9):

QS = 42 where is x or what is x ?

jhonyy9 (jhonyy9):

ok. i see it now and you are right sure

OpenStudy (anonymous):

yes, and sorry for the wait laptop messing up

jhonyy9 (jhonyy9):

ok. so can you prove it substituting the 13 in place of x inside this equation and you need to get 42

OpenStudy (anonymous):

yes, but i am also suppose to write the reasons... which i am not good at...

jhonyy9 (jhonyy9):

but i dont understand it what you need to reason it 13+3 +2*13 = 42 16+26 = 42 42 =42 in this way this is proven x=13 is the right answer sure

OpenStudy (anonymous):

to justify each step is what it says.....

jhonyy9 (jhonyy9):

QR = x+3 RS =2x are these given ?

jhonyy9 (jhonyy9):

and so x+3 + 2x =42 are given too ?

OpenStudy (anonymous):

yes i forgot to add that... sorry the two on top is in the timeline thingy.. but the equations is not